OpenStudy (veronicamorgan2):

What is the 10th term of the sequence represented by an = -3n + (-1)n? -8

Hayhayz (hayhayz):

Hi Welcome to OpenStudy! You will get help faster if you post your question in the correct section :) You can find subjects by clicking on "OpenStudy Feedback" !

OpenStudy (veronicamorgan2):

What is the 10th term of the sequence represented by an = -3n + (-1)n? -8

Vocaloid (vocaloid):

post your question in the math section please http://openstudy.com/study#/groups/Mathematics

OpenStudy (mathmale):

Your expression, an = -3n + (-1)n, is ambiguous in two places. If you take pains to present it properly, you might get faster and better responses from would-be helpers. an: Your intention was to mention "the nth term of an arithmetic sequence." Please type that as a_1 ("a sub n"), or, if you want to be both clear and fancy, type it in the Equation Editor (below) as \[a _{n}\]

OpenStudy (mathmale):

Secondly, by "(-1)n," I think you meant \[(-1)^n\]which is "the nth power of -1."

OpenStudy (mathmale):

If you did mean " a_n = -3n + (-1)^n, then: a_1=-3(1)+(-1)^1=-3-1=-4 Try figuring out a_2, a_3, etc. Finally find a_10.